if x is a positive real number, show thata x+(4/x)>=4. When does equality hold?
Answer by venugopalramana(3286)

if x is a positive real number, show thata x+(4/x)>=4. When does equality hold?
LET Y=X+4/X-4
Y=(X^2+4-4X)/X
=(X-2)^/X
N.R. BEING A PERFECT SQUARE IS ALWAYS POSITIVE.X IS GIVEN TO BE POSITIVE.HENCE Y IS POSITIVE.ITS MINIMUM VALUE WOULD BE ZERO WHEN X=2
THAT IS X+4/X=MINIMUM VALUE OF 4 WHEN X=2